OpenStudy (blockcolder):

Serious question: Prove that for any positive integer n there exists a positive multiple of n that contain only the digits 7 and 0. (in decimal system, of course.) I know I'm supposed to use some sort of the pigeonhole principle, but I have no idea what the pigeons and pigeonholes are. Any hints?

OpenStudy (experimentx):

1*7 = 7 2*35 = 70 3*259=777 4*175=700 5*14 = 70 6*1295 = 7770 7*10 = 70 .. .. looks like we really can construct it.

OpenStudy (thomas5267):

This is the pigeonhole principle, but I have no idea how to use this to prove the question. http://en.wikipedia.org/wiki/Pigeonhole_principle

OpenStudy (experimentx):

I hate to do it ... but there is answer in google ... take a peek if you like!!! :(

OpenStudy (anonymous):

Looked at the solution on google, and it makes sense, but I would never have figured that out myself.

OpenStudy (experimentx):

me ... not in next 100 years!!!

OpenStudy (blockcolder):

My solution was close, but I didn't realize I could just subtract the 2 numbers in the same pigeonhole. I thought I had to prove that one of pigeons should lie in the pigeonhole labeled 0. Anyway, thanks for telling me about the internet solution. This is what I get for enrolling in a problem-solving course. Closing in 3, 2, 1...